What is (2x-2) + (-2x+18) + (7x-27) = 24
Solve for x.

Answer by ikleyn(52781)

.
1)  In the left side, combine like terms with x:

    2x - 2x + 7x = 7x.



2)  In the left side, combine constant terms

    -2 + 18 - 27 = -11.



3)  Thus your equation takes the form 

    7x - 11 = 24.



4)  Add 11 to both sides. You will get

    7x = 24 + 11 = 35.



5)  The last step is to divide both sides by 7. When you do it, you will get the solution
 
    x = 35/7 = 5.


ANSWER.  x= 5.

Solved.